sacar ib, ic, ie, vcb, vbe, vbb,
7.2k -20 C2 Ovo C1 01 vin 0. B=45 개 100k -9V
Answer #1

Sal DC Analysis ал с рамо, о5у ором ск. о 1.2k Nь, le в – 45 Ve, + -N 1 be (ook 1 . b - V Ve = 0.6Vo), іс - вi = Чsi, ie = 1b6=-8+ Vibe Vbb =-g+0.6 Ub6-gouvalt Nce= -2-7.2kic = -2 -7.2K* 3.73 m Ve=-29.216 valt Vect Wcb = -29.216-(-8.4) Vch = 20.816